OpenStudy (anonymous):

I don't understand how to exactly do this.. APR 19.1% compounded monthly Grace Period 30 days Minimum Payment 3.5% of the remaining balance or $10, whichever is higher Patti charged $425 to her credit card last month. What is the first minimum payment she must make?

OpenStudy (amistre64):

what is 3.5% of 425? is it bigger or smaller than 10?

OpenStudy (anonymous):

yes, it came out to 15.75

OpenStudy (amistre64):

then thats the minimum payment if calculated correctly :)

OpenStudy (anonymous):

but that answer isn't one of choices

OpenStudy (amistre64):

is 14.88?

OpenStudy (anonymous):

yes, I put that at my answer because it was close to what I got but I don't understand how I calculated it wrong.

OpenStudy (anonymous):

ohh. I multiplied by .35 instead of .035

OpenStudy (amistre64):

if there had been no grace period: 425(1+.191/12) = 431.76 3.5% of 431.76 = 15.12 ----------------------------------- BUT, there is a grace period so the purchase from last month is not compounded, so just take: 3.5% of 425 = 14.88

OpenStudy (anonymous):

Yeah, I keep forgetting to move the decimal two spots. but can you help me with on more problem?

OpenStudy (amistre64):

on more, maybe

OpenStudy (anonymous):

Nash uses a credit card with a 17.8% APR, compounded monthly, to pay for a cruise totaling $1,789.17. He can pay $650 per month on the card. What will the total cost of this purchase be? Isn't that the same equation you used up there?

OpenStudy (amistre64):

its related, but this takes on a rather new form

OpenStudy (amistre64):

let B = 1789.17, the balance to pay off; let k = the compounding jargon (1+.178/12), and let P be the payment M0 = B M1 = Bk - P M2 = Bk^2 -Pk - P M3 = Bk^3 -Pk^2 -Pk - P M4 = Bk^4 -Pk^3 -Pk^2 -Pk - P .... Mn = Bk^n -P(1+k+k^2+...+k^(n-1)) ^^^^^^^^^^^^^^^^^ geometrim sum in k Mn = Bk^n -P(1-k^n)/(1-k) they key is in solvinig this setup for n

OpenStudy (amistre64):

let Mn = 0 be the amount when its paid off 0 = Bk^n -P(1-k^n)/(1-k) Bk^n = P(1-k^n)/(1-k) B(1-k)/P k^n = 1- k^n B(1-k)/P = (1- k^n)/k^n B(1-k)/P = 1/k^n - 1 (B(1-k)/P) + 1 = 1/k^n 1/((B(1-k)/P)+1) = k^n ln(1/((B(1-k)/P)+1)) = n ln(k) ln(1/((B(1-k)/P)+1))/ln(k) = n thats was fun ....

OpenStudy (amistre64):

ln(1/((B(1-k)/P)+1))/ln(k) B = 1789.17 k = 1+.178/12 P = 650 ln(1/((1789.17(1-k)/650)+1))/ln(k), k = 1+.178/12 http://www.wolframalpha.com/input/?i=ln%281%2F%28%281789.17%281-k%29%2F650%29%2B1%29%29%2Fln%28k%29%2C+k+%3D+1%2B.178%2F12 id say about 3 months to pay off: or simply 650 * n, about 1840.25

OpenStudy (anonymous):

that was so much work, but I got it now.

OpenStudy (amistre64):

we could of course have just gone with a couple of iterations .. 1789.17(1+.178/12) - 650 (1789.17(1+.178/12) - 650)(1+.178/12)-650 ((1789.17(1+.178/12) - 650)(1+.178/12)-650)(1+.178/12); remaining balance 650+650+remaining balance

OpenStudy (amistre64):

1840.90 if we go that route

OpenStudy (dumbcow):

just to add...this takes 30 sec with a financial calculator depending on level of class you are taking, you may not need to know all the math behind the calculations

OpenStudy (amistre64):

from the blank stares that i see from the accounting majors as they gaze upon their financial calculators in fear and awe ... im pretty sure those things eat your brain :)